OpenStudy (anonymous):

Use the following function to complete the question below below: f(x) = 2(x-6)^2+4 Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

Is this the entire question or is there more?

OpenStudy (anonymous):

@BrooklynRoyals

OpenStudy (anonymous):

Oh, I read that wrong. @GodGirl360, this is the entire question.

OpenStudy (anonymous):

@cwrw238, can you help me?

OpenStudy (cwrw238):

yes this is quite straight forward first find the value of f(x) when x = 7 and x = 4

OpenStudy (anonymous):

Oh, okay! So I just plug those numbers in and solve for them?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Ok, let me try working it out.

OpenStudy (cwrw238):

eg f(7) = 2(7-6)^2 = 4

OpenStudy (cwrw238):

sorry that is + 4

OpenStudy (anonymous):

f(x) = 2(x-6)^2+4 f(4) = 2(4-6)^2-4 4 = 2(-2)^2-4 4 = 2(4) -4 4 = 8 - 4 4 =4 Is that correct?

OpenStudy (cwrw238):

nope f(4) = 2(4-6)^2 + 4 = 2(-2)^2 + 4 = 2*4 + 4 = 8 +4 = 12

OpenStudy (anonymous):

Oh, right, it's +4

OpenStudy (anonymous):

Let me try the other one

OpenStudy (cwrw238):

you are not solving an equation you are just finding the value of f(x) when x = 4

OpenStudy (cwrw238):

oh - i confused you with the = !

OpenStudy (cwrw238):

that was a typo

OpenStudy (anonymous):

It wont let me post my reply, one second

OpenStudy (anonymous):

f(7) = 2(7-6)^2 + 4 7 = 2(1)^2 + 4 7 = 2 + 4 7 = 6 Is this one right?

OpenStudy (cwrw238):

6 is correct i'm not sure why you are putting the 7 there ...

OpenStudy (anonymous):

Sorry, I forgot that it's solving for the value of f(x)

OpenStudy (cwrw238):

anyway the average rate of change will be [f(7) - f(4)] / (7 - 4)

OpenStudy (anonymous):

So it would be (6 - 12)/(7-4)?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Okay, so would I solve it like this?

OpenStudy (anonymous):

Actually, I forgot how to divide this. how do I do that again?

OpenStudy (cwrw238):

you work out the subtractions first

OpenStudy (anonymous):

Oh, okay

OpenStudy (cwrw238):